# Style-Adaptive Detection Transformer for Single-Source Domain Generalized Object Detection

### v1

@@ IV-E Visualization and Analysis
Feature visualization. We employ the t-distributed Stochastic Neighbor Embedding (t-SNE) method[49] to analyze the effectiveness of the proposed DSA and OCL module. First, we separately show the style distribution differences among various domains for the vanilla DINO and SA-DETR, as illustrated in Fig. 3. It can be observed that the style distributions across various domains are well-separated in the vanilla DINO. Conversely, our DSA projects the style representation of the target domain into the source domain, thereby entangling the style distributions across various domains. Through this dynamic adaptation to unseen scenarios, our DSA can more flexibly respond to diverse environmental changes, thereby effectively enhancing the detector’s generalization capability.
@@
Detection results. We present the visualization results of SA-DETR across all scenarios, alongside those of vanilla DINO[39], Clip-gap[5], and Groundtruth, as shown in Fig. 6. Compared to CNN-based methods, DETR-based detectors demonstrate greater accuracy in detecting small and weakly featured objects, which can be attributed to the transformer architecture’s effectiveness in capturing global structural information. Building on this, SA-DETR demonstrates more accurate classification results and higher recall rates when handling unseen scenarios, indicating that our proposed approach effectively enhances the detector’s generalization capability. The visual results align with the numerical evaluations, indicating that our method exhibits excellent performance and generalization capability in both the source domain and various unseen target domains.
@@ V Conclusion
This paper introduces the Style-Adaptive DEtection TRansformer (SA-DETR), a strong DETR-based detector designed for single-source domain generalization (SDG) in object detection. To the best of our knowledge, this is the first work that explores the benefit of a DETR-based detector for the SDG task. The SA-DETR adopts DINO as the base detector and incorporates a domain style adapter and an object-aware contrastive learning module. Firstly, the proposed adapter projects the style representation from the unseen target domain into the training domain, enabling dynamic style adaptation across various unseen scenarios. This establishes a new avenue that utilizes known data to represent unseen ones, thereby successfully adapting to unseen scenarios with significant differences. Then, the proposed module uses carefully designed object-aware gating masks to constrain the scope of contrastive learning in both spatial position and semantic category, forcing the detector to extract domain-invariant features of objects to further enhance the detector’s generalization capability. Extensive experiments across five different weather condition scenarios show that SA-DETR achieves superior performance for SDG in object detection. In future work, we plan to explore methods for domain adaptation during the test-time stage or under conditions of small sample sizes.

### v2

@@ IV-E Visualization and Analysis
IV-E1 Feature visualization
We employ the t-distributed Stochastic Neighbor Embedding (t-SNE) method[62] to analyze the effectiveness of the proposed DSA and OCL module. First, we separately show the style distribution differences among various domains for the vanilla DINO and SA-DETR, as illustrated in Fig. 3. It can be observed that the style distributions across various domains are well-separated in the vanilla DINO. Conversely, our ODS-Adapter projects the style representation of the target domain into the source domain, thereby entangling the style distributions across various domains. Through this dynamic adaptation to unseen scenarios, our adapter can more flexibly respond to diverse environmental changes, thereby effectively enhancing the detector’s generalization capability.
@@ IV-E1 Feature visualization
IV-E2 Detection results
We present the visualization results of SA-DETR across all scenarios, alongside those of vanilla DINO[50], Clip-gap[5], and Groundtruth, as shown in Fig. 6. Compared to CNN-based methods, DETR-based detectors demonstrate greater accuracy in detecting small and weakly featured objects, which can be attributed to the transformer architecture’s effectiveness in capturing global structural information. Building on this, SA-DETR demonstrates more accurate classification results and higher recall rates when handling unseen scenarios, indicating that our proposed approach effectively enhances the detector’s generalization capability. The visual results align with the numerical evaluations, indicating that our method exhibits excellent performance and generalization capability in both the source domain and various unseen target domains.
@@ V Conclusion
This paper introduces the Style-Adaptive DEtection TRansformer (SA-DETR), an effective and efficient DETR-based detector designed for single-source domain generalization (SDG) in object detection. To the best of our knowledge, this is the first work that explores a DETR-based detector for the SDG task. The SA-DETR adopts DINO as the base detector and incorporates an online domain style adapter and an object-aware contrastive learning module. Firstly, the adapter projects the style representation from the target domain onto the source domain, enabling dynamic style adaptation across various unseen scenarios. This adapter constructs a dynamic memory bank that self-organizes into diverse style bases and continuously absorbs statistics from unseen domains under a test-time adaptation framework, facilitating rapid and accurate style adaptation to previously unseen scenarios. Then, the proposed module uses annotations to constrain the scope of contrastive learning in both position and category, encouraging the detector to extract instance-level domain-invariant features to further enhance the detector’s generalization capabilities. Extensive experiments across five distinct scenarios show that SA-DETR achieves superior performance for single-source domain generalization in object detection. In future work, we plan to explore methods for domain adaptation under conditions of small sample sizes.